OpenStudy (anonymous):

14^x+1=36? exponential equations round to the nearest ten thousandth

OpenStudy (anonymous):

Use Log to both sides

OpenStudy (anonymous):

log 14^x+1 = log 36

OpenStudy (anonymous):

It will be (x+1) Log 14=Log 36 (x+1)=Log36/Log14 Can you complete ?

OpenStudy (anonymous):

after you apply the power property of logarithms right? and im not sure

OpenStudy (anonymous):

@kaytiey : Is that your question \[\Large 14^{x+1}=36 \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

In that case ,Do as i did in the pervious steps ; (x+1)=Log36/Log14 x =(Log36/Log14)-1= ?

OpenStudy (anonymous):

previous*

OpenStudy (anonymous):

so .3578 and round it o the nearest ten thousandth 0.358 or what? i dont remember my places after the decimal point

OpenStudy (anonymous):

To the nearest thousandth will be 0.358

OpenStudy (anonymous):

to the nearest ten thousandth

OpenStudy (anonymous):

*i need

OpenStudy (anonymous):

To the nearest ten thousandth ,will be 0.3580 -.

OpenStudy (anonymous):

ok ay the whole number that showed on my calculator is .357878398 so i only round the fourth place after the decimalcorrect?

OpenStudy (anonymous):

Correct :!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Yw